# ListT5: Listwise Reranking with Fusion-in-Decoder Improves Zero-shot Retrieval

### v1

@@ C.2 Modification to Accept ListT5 ($\mathbf{r}$ = 2)
Fig. 7 illustrates and compare the inference scenario for ListT5 ($r$ = 1) and ListT5 ($r$ = 2). Since the ($r$=2) variant outputs 2 different passage indices, the total number of candidate passages we need to compute is multipled by 2 for the next level of a tree. For example, if we use ListT5 ($r$ = 1) with $m$ = 5 at the initial iteration, after computation of the bottom-level, we would get a total of 20 (100 / 5) filtered candidates which will go to the next round for comparison. In contrast, if we use ListT5 ($r$ = 2) with $m$ = 5, we would get a total of 40 (100/5 *** 2) candidates to compute at the next round. However, it is important to note that we only use the inference method of ($r$ = 1) for the root computation, since we always outputs the top-1 candidates for every iteration.
@@ C.3 Details on Random Replacement
The basic unit of ListT5 always accepts exactly $m$ passages as the input. However, there are cases when the leftover passages are not a multiple of $m$. For example, as we proceed to the upper level on the tournament tree, there will be a total of 100 $\mapsto$ 20 $\mapsto$ 4 passages left if we rank 100 passages using ListT5 ($r$=1) with $m$=5. In this case, we would need one more dummy passage to make 4 candidate passages to 5. For reproducibility, we select a dummy passage in a deterministic way by considering the first passage from the inital candidate passage pool. If that passage is already selected as top-$k$ indices during the previous round or we have duplicates inside the input window $m$, we proceed to the next index until the conditions are met.
@@ C.4 Example Scenario using Output Caching
Table 8: NDCG@10 results for reranking on top of COCO-DR large Top-100 first-stage retrieval results.
[t] We also conduct reranking using the Top-100 results of COCO-DR-large Yu et al. (2022), and compare reranked results with MonoT5 Nogueira et al. (2020), RankT5 Zhuang et al. (2022), and ListT5. For this experiment, we could only utilize the evaluation subset that is uploaded (fully open) in the BEIR repository. [^7] Thus, results for BioASQ, Signal-1M, TREC-NEWS, Robust04 are not computed. For simplicity, the input token length is fixed to 512 for first-stage retrieval with COCO-DR. From the results at Tab 8, ListT5 achieves an average of +1.5 point gain for reranking on COCO-DR Top-100, additionally showing its applicability and effectiveness on both lexical-based (BM25) and neural-based first-stage retrieval modules.
@@ Appendix E MRR scores
Table 9: Reporting Mean Reciprocal Rank (MRR)@10 scores for the main table.
@@ Appendix G Details on Evaluation Dataset
We download and use the officially released checkpoints for the baseline model with the huggingface identifier of castorini/monot5-base-msmarco-10k, monot5-3b-msmarco-10k for MonoT5 Nogueira et al. (2020), castorini/duot5-base-msmarco for DuoT5 Pradeep et al. (2021), castorini/rank_vicuna_7b_v1 for RankVicuna Pradeep et al. (2023a), and castorini/rank_zephyr_7b_v1_full for RankZephyr Pradeep et al. (2023b). For RankGPTSun et al. (2023b) using GPT-3.5 or GPT-4, we evaluate them using the OpenAI OpenAI (2022) API. For RankT5, we download the released T5X checkpoint from the official repository [^8] https://github.com/google-research/google-research/tree/master/rankt5 and convert it into a PyTorch checkpoint using the HuggingFace’s conversion script999convert_t5x_checkpoint_to_pytorch.py. We evaluate DuoT5 on the same setup to the paper’s final experiments - we rerank 100 passages using MonoT5, select top 50 passages to be run on DuoT5, and aggregate relevancy scores of individual documents by the SYM-SUM method.
@@ G.3 Other Replication Details
We run the same evaluation setup with the BEIR library to do evaluation. For example, we append title with a space along with the text(passage) part and treat as full passage. For evaluating reranked results for both BM25 Top100 and COCO-DR Top100 candidate passages, we remove duplicates before evaluation [^13] For example, quora had 1 (out of 10000) duplicates, so the NDCG@10 for BM25 may differ from the BEIR paper. For CQADupStack, we follow the same procedure and aggregated all sub-domains ranging from android to wordpress to make 13,145 test queries, and report the average performance.
@@ H.3 Detailed analysis about pairwise methods
H.4 Experiments on the consistency of LLMs.
GPT-3.5-turbo-1106 ListT5 -base
Trial 1 Trial 2 Trial 3 Avg.
(1) Accuracy when the gold passage is at index #:
1 81.6 79.6 81.6 81.0 93.9
2 63.3 63.3 61.2 62.6 87.8
3 75.5 75.5 75.5 75.5 83.7
4 67.3 63.3 67.3 66.0 85.7
5 61.2 63.3 65.3 63.3 81.6
std 7.68 7.1 7.4 7.4 4.2
(2) Agreement ratio (%) within index change of positive
points to
same passage
55.1 55.1 55.1 55.1 83.7
other 44.9 44.9 44.9 44.9 16.3
Table 10: Measuring the LLM consistency on TREC-COVID.
@@
Table 11: Example of ListT5, MonoT5, RankT5 retrieval result on Arguana dataset. Relevant topics are highlighted with red, non-relevant topics in blue, and opinions are bolded.
@@ Appendix I Qualitative Analysis
sliding window, stride=1 1 + $\lceil$(100-5)/1$\rceil$ = 96 96 $\times$ $\lceil$10/4$\rceil$ = 288
sliding window, stride=2 1 + $\lceil$(100-5)/2$\rceil$ = 49 49 $\times$ $\lceil$10/3$\rceil$ = 196
sliding window, stride=3 1 + $\lceil$(100-5)/3$\rceil$ = 33 33 $\times$ $\lceil$10/2$\rceil$ = 165
sliding window, stride=4 1 + $\lceil$(100-5)/4$\rceil$ = 25 25 $\times$ $\lceil$10/1$\rceil$ = 250
@@
Table 12: Number of forward passes to rerank top-k candidates from 100 candidate passages per one query, where window size $w$=5. In the case of reranking top-10 passages, tournament sort requires much more fewer number of forward passes.
@@ 0 T5-base pointwise MonoT5 1x 1x
Table 13: FLOPs (In a multiple of FLOPs of MonoT5-base) on the choice of architecture and method, on TREC-DL19. For the sliding window approach, we would need a total of 4 multiple passes for stride = 3 and 5 passes for stride = 2 (Explained at Tab. 12) to rerank Top-10 candidates.
@@
Table 14: Comparison of FLOPs and performance on ListT5 with the S.W (Sliding Window) approach and T.S (Tournament Sort) with different hyperparameters. ListT5 ($r$=2) performs the best, with lower FLOPs than the sliding window variants on both setup of reranking top-10 and top-1 candidates.
[t]
Appendix J Detailed Comparison with the Sliding Window approach with respect to Tournament Sort.
In this section, we analyze the efficiency of tournament sort v.s. the sliding window approach.
@@
We rerank top-10 passages from BM25 top100 candidate passages, for 43 queries of TREC-DL19. The latency and FLOPs for tournament sort variants are computed using our ListT5 evaluation code, and the efficiency for the sliding window approach (without the FiD architecture) is computed using the official repository for RankVicuna and RankZephyr. [^16] https://github.com/castorini/rank_llm The FiD variant of the sliding window approach are computed using our ListT5 evaluation code. We append the deepspeed FlopsProfiler for FLOPs measurement, which is the same as the ones that was appended for ListT5. For a fair comparison, we also measure the FLOPs of the sliding approach on the T5-3b model171717we replace the model code from castorini/rank_vicuna_7b_v1 to lmsys/fastchat-t5-3b-v1.0 and measure FLOPs. and the T5-base FiD architecture using the ListT5 code. Example commands used are as follows:
@@
We calculate the number of required forward passes needed for variants of tournament sort and the sliding window approach at Table  12. Given a window size of 5, tournament sort is much more efficient to rerank top-10 passages, requiring fewer number of forward passes. This is because, unlike tournament sort with output caching, the sliding window approach requires re-evaluation over the entire input sequence, depending on the window size, stride, and the number of top-k passages to rerank. Detailed explanation of how we calculated the numbers from the table is below:
  • After one pass of a sliding window of =5 and stride of 4, we discard 4 passages and only carry (5-4)=1 previous passage to the next step as we move the window. Therefore, we would only be able to correctly order top-1 passages, since the top-2 passage cannot be moved.
  • Therefore, to rank top 10 candidates, we have to iterate through at least $\lceil$10 / (5-4$\rceil$ = 10 times, which would result in a total of 240 forward passes.
@@
Performance Comparison. Comparison of FLOPs and performance on ListT5 with the S.W(Sliding Window) approach and T.S(Tournament Sort) with different hyperparameters. ListT5 ($r$=2) performs the best, with lower FLOPs than the sliding window variants on both setup of reranking top-10 and top-1 candidates.

### v3

@@ C.2 Modification to Accept ListT5 ($\mathbf{r}$ = 2)
Fig. 7 illustrates and compare the inference scenario for ListT5 ($r$ = 1) and ListT5 ($r$ = 2). Since the ($r$=2) variant outputs 2 different passage indices, the total number of candidate passages we need to compute is multipled by 2 for the next level of a tree. For example, if we use ListT5 ($r$ = 1) with $m$ = 5 at the initial iteration, after computation of the bottom-level, we would get a total of 20 (100 / 5) filtered candidates which will go to the next round for comparison. In contrast, if we use ListT5 ($r$ = 2) with $m$ = 5, we would get a total of 40 (100/5 * 2) candidates to compute at the next round. However, it is important to note that we only use the inference method of ($r$ = 1) for the root computation, since we always outputs the top-1 candidates for every iteration.
@@ C.3 Details on Random Replacement
Random replace index number:
Dataset
21
20 19 18 17 16 15 6
signal 33.5 33.3 33.2 33.1 33.3 33.3 33.3 33.5
trec-covid 78.3 78.2 78.2 78.4 78.1 78.1 78.3 78.3
news 48.5 48.4 48.4 48.4 48.3 48.4 48.4 48.4
scifact 74.1 74.0 74.0 74.0 74.0 74.1 74.2 74.0
nfcorpus 35.6 35.7 35.7 35.7 35.7 35.7 35.7 35.8
fiqa 39.6 39.6 39.7 39.7 39.6 39.6 39.6 39.7
bioasq 56.4 56.6 56.5 56.5 56.4 56.5 56.5 56.5
touche 33.4 33.7 33.4 33.5 33.6 33.8 33.8 33.6
dbpedia 43.7 43.6 43.8 43.7 43.8 43.7 43.7 43.8
robust04 52.1 52.1 52.0 52.1 52.0 52.1 52.1 51.9
scidocs 17.6 17.6 17.6 17.5 17.6 17.5 17.6 17.5
arguana 48.9 49.0 48.8 48.8 48.9 49.0 48.9 48.9
climate-fever 24.0 24.0 24.0 24.0 24.0 24.0 24.0 24.0
dl19 71.8 71.9 71.8 71.7 71.7 71.8 71.7 71.8
dl20 68.1 68.2 68.1 68.1 68.1 68.1 68.3 68.5
Average 48.4 48.4 48.3 48.3 48.3 48.4 48.4 48.4
Table 9: NDCG@10 results on differing the index of random assignment, with explanations on Section C.3.
The basic unit of ListT5 always accepts exactly $m$ passages as the input. However, there are cases when the leftover passages are not a multiple of $m$. For example, as we proceed to the upper level on the tournament tree, there will be a total of 100 $\mapsto$ 20 $\mapsto$ 4 passages left if we rank 100 passages using ListT5 ($r$=1) with $m$=5. In this case, we would need one more dummy passage to make 4 candidate passages to 5. For reproducibility, we select a dummy passage in a deterministic way by considering the first passage from the initial candidate passage pool. If that passage is already selected as top-$k$ indices during the previous round or we have duplicates inside the input window $m$, we proceed to the next index until the conditions are met. To show that the assignment of the random index is not statistically significant, we have conducted additional experiments on the subset of BEIR with differing the random replace index number. The results are presented at Table. 9. From the table, we can see that the output is robust to the assignment of random replace index number.
@@ C.4 Example Scenario using Output Caching
Table 10: NDCG@10 results for reranking on top of COCO-DR large Top-100 first-stage retrieval results (Sec. 10).
We also conduct reranking using the Top-100 results of COCO-DR-large Yu et al. (2022), and compare reranked results with MonoT5 Nogueira et al. (2020), RankT5 Zhuang et al. (2022), and ListT5. For this experiment, we could only utilize the evaluation subset that is uploaded (fully open) in the BEIR repository. [^9] Thus, results for BioASQ, Signal-1M, TREC-NEWS, Robust04 are not computed. For simplicity, the input token length is fixed to 512 for first-stage retrieval with COCO-DR. From the results at Tab 10, ListT5 achieves an average of +1.5 point gain for reranking on COCO-DR Top-100, additionally showing its applicability and effectiveness on both lexical-based (BM25) and neural-based first-stage retrieval modules.
@@ Appendix E MRR scores
Table 11: Reporting Mean Reciprocal Rank (MRR)@10 scores for initial ranking of BM25 top-100 and COCO-DR large top-100. (Sec. E)
@@ Appendix G Details on Evaluation Dataset
We download and use the officially released checkpoints for the baseline model with the huggingface identifier of castorini/monot5-base-msmarco-10k, monot5-3b-msmarco-10k for MonoT5 Nogueira et al. (2020), castorini/duot5-base-msmarco for DuoT5 Pradeep et al. (2021), castorini/rank_vicuna_7b_v1 for RankVicuna Pradeep et al. (2023a), and castorini/rank_zephyr_7b_v1_full for RankZephyr Pradeep et al. (2023b). For RankGPTSun et al. (2023b) using GPT-3.5 or GPT-4, we evaluate them using the OpenAI OpenAI (2022) API. For RankT5, we download the released T5X checkpoint from the official repository [^10] https://github.com/google-research/google-research/tree/master/rankt5 and convert it into a PyTorch checkpoint using the HuggingFace’s conversion script111111convert_t5x_checkpoint_to_pytorch.py. We evaluate DuoT5 on the same setup to the paper’s final experiments - we rerank 100 passages using MonoT5, select top 50 passages to be run on DuoT5, and aggregate relevancy scores of individual documents by the SYM-SUM method.
@@ G.3 Other Replication Details
Output validation module Upon experiments, we do not use constrained decoding or validation module to ensure the generation of valid permutations of 5. Each digital number is represented as simple integers (The decoded output looks something like "1 2 5 4 3" or "2 3 1 5 4"), with an exception catching module such that if the parsing fail, we go back to the original ordering. Even without any validation module, since ListT5 is fine-tuned to output valid indexes for 20000 steps, if the input follows the correct format, the model almost always outputs valid index. Nevertheless, we think that adding a constrained decoding module would be beneficial.
Evaluation We run the same evaluation setup with the BEIR library to do evaluation. For example, we append title with a space along with the text(passage) part and treat as full passage. For evaluating reranked results for both BM25 Top100 and COCO-DR Top100 candidate passages, we remove duplicates before evaluation [^15] For example, quora had 1 (out of 10000) duplicates, so the NDCG@10 for BM25 may differ from the BEIR paper. For CQADupStack, we follow the same procedure and aggregated all sub-domains ranging from android to wordpress to make 13,145 test queries, and report the average performance.
@@ H.3 Detailed analysis about pairwise methods
Table 12: Example of ListT5, MonoT5, RankT5 retrieval result on Arguana dataset. Relevant topics are highlighted with red, non-relevant topics in blue, and opinions are bolded. (Sec. I)
H.4 Experiments on the consistency of LLMs.
GPT-3.5-turbo-1106 ListT5 -base
Trial 1 Trial 2 Trial 3 Avg.
(1) Accuracy when the gold passage is at index #:
1 81.6 79.6 81.6 81.0 93.9
2 63.3 63.3 61.2 62.6 87.8
3 75.5 75.5 75.5 75.5 83.7
4 67.3 63.3 67.3 66.0 85.7
5 61.2 63.3 65.3 63.3 81.6
std 7.68 7.1 7.4 7.4 4.2
(2) Agreement ratio (%) within index change of positive
points to
same passage
55.1 55.1 55.1 55.1 83.7
other 44.9 44.9 44.9 44.9 16.3
Table 13: Measuring the LLM consistency on TREC-COVID. (Sec. H.4)
@@ Appendix I Qualitative Analysis
sliding window, stride=1 1 + $\lceil$(100-5)/1$\rceil$ = 96 96 $\times$ $\lceil$10/4$\rceil$ = 288 -> 280
sliding window, stride=2 1 + $\lceil$(100-5)/2$\rceil$ = 49 49 $\times$ $\lceil$10/3$\rceil$ = 196 -> 191
sliding window, stride=3 1 + $\lceil$(100-5)/3$\rceil$ = 33 33 $\times$ $\lceil$10/2$\rceil$ = 165 -> 162
sliding window, stride=4 1 + $\lceil$(100-5)/4$\rceil$ = 25 25 $\times$ $\lceil$10/1$\rceil$ = 250 -> 248
@@
Table 14: Number of forward passes to rerank top-k candidates from 100 candidate passages per one query, where window size $w$=5. In the case of reranking top-10 passages, tournament sort requires much more fewer number of forward passes. (Sec. J.2)
@@ 0 T5-base pointwise MonoT5 1x 1x
Table 15: FLOPs (In a multiple of FLOPs of MonoT5-base) on the choice of architecture and method, on TREC-DL19. For the sliding window approach, we would need a total of 4 multiple passes for stride = 3 and 5 passes for stride = 2 (Explained at Tab. 14) to rerank Top-10 candidates. (Sec. J.3)
@@
Table 16: Comparison of FLOPs and performance on ListT5 with the S.W (Sliding Window) approach and T.S (Tournament Sort) with different hyperparameters. ListT5 ($r$=2) performs the best, with lower FLOPs than the sliding window variants on both setup of reranking top-10 and top-1 candidates. (Sec. J.4)
Appendix J Measuring efficiency - Sliding window v.s. Tournament Sort.
Figure 8: Overview comparison of sliding window v.s. Tournament sort for listwise reranking. (Sec. 8)
We briefly describe the difference of tournament sort and the sliding window approach in Figure. 8. We conduct additional experiments to compare the efficiency of tournament sort and the sliding window.
@@
We rerank top-10 passages from BM25 top100 candidate passages, for 43 queries of TREC-DL19. The latency and FLOPs for tournament sort variants are computed using our ListT5 evaluation code, and the efficiency for the sliding window approach (without the FiD architecture) is computed using the official repository for RankVicuna and RankZephyr. [^18] https://github.com/castorini/rank_llm The FiD variant of the sliding window approach are computed using our ListT5 evaluation code. We append the deepspeed FlopsProfiler for FLOPs measurement, which is the same as the ones that was appended for ListT5. For a fair comparison, we also measure the FLOPs of the sliding approach on the T5-3b model191919we replace the model code from castorini/rank_vicuna_7b_v1 to lmsys/fastchat-t5-3b-v1.0 and measure FLOPs. and the T5-base FiD architecture using the ListT5 code. Example commands used are as follows:
@@
We calculate the number of required forward passes needed for variants of tournament sort and the sliding window approach at Table  14. Given a window size of 5, tournament sort is much more efficient to rerank top-10 passages, requiring fewer number of forward passes. This is because, unlike tournament sort with output caching, the sliding window approach requires re-evaluation over the entire input sequence, depending on the window size, stride, and the number of top-k passages to rerank. Detailed explanation of how we calculated the numbers from the table is below:
  • After one pass of a sliding window of size 5 and stride of 4, we discard 4 passages and only carry (5-4)=1 previous passage to the next step as we move the window. Therefore, we would only be able to correctly order top-1 passages, since the top-2 passage cannot be moved.
  • Therefore, to rank top 10 candidates, we have to iterate through at least $\lceil$10 / (5-4$\rceil$ = 10 times, which would result in a total of about 250 forward passes.
  • For ranking top-6 with a sliding window of size 5 and stride 4, some forward pass can be saved in the 6th slide because top-5 have already been ordered in the first 5 slides. Therefore, a corrected precise calculation (noted as -> in the table) gives 248.
@@
J.4 Performance Comparison.
Comparison of FLOPs and performance on ListT5 with the S.W(Sliding Window) approach and T.S(Tournament Sort) with different hyperparameters. ListT5 ($r$=2) performs the best, with lower FLOPs than the sliding window variants on both setup of reranking top-10 and top-1 candidates.
Appendix K Additional Experiments on Tournament Sort.
Figure 9: Illustration for the tournament sort to rank top-10 passages among 100 candidates for $r$=10 and $m$=20. It takes 9 forward passes with unwanted wins. (Sec. K.1)
# of stride == value of r
10 5 1
Sliding w. 9 x 1 iter. = 9 17 x 2 iter. = 34 81 x 10 iter. = 810
Tournament S. 9 x 1 iter. = 9 7 x 2 iter. = 14 6 x 10 iter. = 60
Table 17: Comparison of the number of forward passes needed to rank top-10 passages, when $m$ = 20, and $n$ = 100. The number is written in the format of {number of forward pass for each iteration} x {number of iterations to assure global ranking}. (Sec. K.1)
K.1 Cases where m > k
In the main paper, we have discussed that the sliding window approach can be much more efficient when $m$ << $k$. However, even when $m$ > $k$, e.g., $m$ = 20 and $k$ = 10, the number of forward passes needed to get top-10 rankings for tournament sort can less than or equal to the sliding window variants, with simple modifications. For example, the number of iterations to rank top-10 with $m$=20 and $r$ = 10 with ListT5 requires 9 forward passes if we take into account all top-$r$ results from one iteration and take into account unwanted wins. Tournament sort with $r$=10 can correctly rank global top-10 candidates in one iteration, as illustrated at Fig. 9. This is the same amount of iterations needed with the sliding window approach with window size of 20 and stride of 10. We also compute the number of forward passes with different value of $s$, or $r$ in Table. 17.
K.2 Tournament Sort with LLMs.
Method dl19 dl20 trec-covid news touche
sliding 68.4 ± 0.4 64.9 ± 1.1 72.6 ± 1.4 46.5 ± 1.0 38.2 ± 0.5
tournament 67.4 ± 0.9 65.8 ± 0.6 76.4 ± 0.4 45.5 ± 1.0 33.1 ± 1.7
Table 18: NDCG@10 on the selected subset of BEIR, on RankGPT-3.5 with different sorting methods. For fair comparison, we used w = 20, s = 10 for the sliding approach, and m = 20, r = 10 for the tournament sort. To compensate for the instability of APIs, all results are run for 3 times. Except for trec-covid and touche, differences are statistically non-significant (p > 0.1). (Sec. 18)
We have also analyzed the performance of RankGPT Sun et al. (2023b), listwise reranking with GPT3.5, with tournament sort. The performance difference of tournament sort with respect to the sliding window approach for w=20, s=10 were not significant, while the number of required forward passes to rank top-10 passages were the same (9) for both variants.